14. An inverse demand function for board games is given by p = 240 - q0.5. Monopoly & Co. currently sells board games for £120, but they are thinking about revising their pricing. Help them with that task by answering the following questions
a. What is the quantity demanded when p = £120?
q =
b. What is the price elasticity of demand at the current price?
Ep=
1
Expert's answer
2019-07-01T09:06:34-0400

a. p = 240 - q^0.5, so q = (240 - p)^2

The quantity demanded when p = £120 is:

q = (240 - 120)^2 = 14400 units.

b. The price elasticity of demand at the current price is:

Ep = -480×120/14400 = -4.